OpenStudy (anonymous):

PLZ HELP WILL GIVE MEDAL!!!!!!! Which graph is the line y=1/3x-2

OpenStudy (anonymous):

just me a sec to post the options

OpenStudy (anonymous):

to find the y intercept substitute y and x with 0. 0 = 1/3(0) - 2 = -2 the y intercept is -2.

imqwerty (imqwerty):

slope = 1/3 y intercept = -2

OpenStudy (meowlover17):

D

OpenStudy (anonymous):

K THANKS

OpenStudy (meowlover17):

Your Answer would be D, Dont forget rise over run.

OpenStudy (anonymous):

yeah it's d

OpenStudy (meowlover17):

Rise 1, run 3, starting at -2

OpenStudy (anonymous):

YEAH it was D

OpenStudy (meowlover17):

:D

OpenStudy (meowlover17):

Awesome